# Imaginary Gauge-steerable Edge Modes In Non-Hermitian Aubry-André-Harper Model

### v1

@@ Abstract
We investigate a non-Hermitian Aubry-André-Harper lattice exhibiting quasiperiodicity, featuring an imaginary gauge field that varies spatially but averages to zero. In the presence of open boundary conditions, this system is precisely mapped, through a nonunitary gauge transformation, to the Hermitian AAH model with balanced hopping terms. The mapping leaves the spectrum unchanged but reshapes each eigenfunction by a realization-dependent random-walk envelope. In a parameter regime where the Hermitian counterpart hosts spectrally isolated in-gap boundary modes, we identify two such modes with sharply different responses to the envelope: one stays anchored at the boundary, while the other is controllable via the gauge, allowing its peak intensity to be relocated solely by altering the gauge setup without modifying the associated eigenenergy. Additionally, we demonstrate that this steerable mode can be preferentially enhanced and generated from an initial bulk wavefunction by introducing mild site-specific amplification at a location determined exclusively from the Hermitian model using the biorthogonal function. These findings offer pathways for both static and dynamic manipulation of spatially adjustable in-gap states in quasiperiodic non-Hermitian lattices.
@@ I Introduction
Non-Hermitian Hamiltonians offer an effective framework for describing driven or open quantum systems characterized by gain and loss, leakage, measurement backaction, or nonreciprocal couplings. These systems support phenomena that have no counterparts in Hermitian regimes, including parity-time-symmetry breaking, exceptional points, and topological structures of complex spectra Bender1998_PRL ; Bender2007_RPP ; Rotter2009_JPA ; Heiss2012_JPA ; ElGanainy2018_NatPhys ; MiriAlu2019_Science ; Ozdemir2019_NatMat ; Ruter2010_NatPhys ; Regensburger2012_Nature ; Hodaei2017_Nature ; Feng2017_NatPhoton ; Bergholtz2021_RMP ; Ashida2020_AdvPhys . In lattice systems, non-Hermiticity disrupts the conventional bulk-boundary correspondence and localization behaviors, thereby facilitating boundary-state accumulation, amplification, and wave manipulation that transcend Hermitian frameworks. NHSE_review_Zhang2022 ; NonHermitianReview_Kawabata2020 .
@@
A hallmark manifestation of non-Hermiticity in lattice systems is the non-Hermitian skin effect (NHSE). Under open boundary conditions (OBC), a macroscopic number of eigenstates become exponentially localized near the boundaries, in stark contrast to the delocalized Bloch waves supported by the same system under periodic boundary conditions (PBC). This pronounced sensitivity to boundary conditions underscores the breakdown of conventional Bloch band theory and has motivated the development of generalized bulk-boundary correspondences based on non-Bloch band theory and point-gap topology YaoWang2018_PRL ; Kunst2018_PRL ; Gong2018_PRX ; YokomizoMurakami2019_PRL ; Kawabata2019_PRX ; Okuma2020_PRL ; Lee2016_PRL . The NHSE and its consequences for boundary spectra and experimentally accessible signatures have been extensively explored Bergholtz2021_RMP ; Ashida2020_AdvPhys ; NHSE_review_Zhang2022 ; NonHermitianReview_Kawabata2020 .
@@
Imaginary gauge fields provide an alternative approach to non-Hermitian localization that does not require uniform nonreciprocity. In the Hatano-Nelson model, a uniform imaginary vector potential alters the localization properties and induces a pronounced sensitivity to boundary conditions HatanoNelson1996_PRL ; HatanoNelson1997_PRB . More recently, Longhi demonstrated that strong skin-like localization can emerge solely from spatial fluctuations in the imaginary gauge field profile, even when the spatial average of the imaginary gauge field vanishes, thereby rendering the lattice globally reciprocal ENHSE_PRL2025 . In the erratic non-Hermitian skin effect (ENHSE), the eigenstates exhibit an effective random-walk envelope, with localization centers determined by the extrema of the corresponding cumulative gauge field. This mechanism is distinct from both the conventional non-Hermitian skin effect, arising from net nonreciprocity, and Anderson localization. Within this framework, a question arises about how boundary-localized in-gap modes respond when such erratic gauge fluctuations are incorporated into quasiperiodic structures.
@@
Quasiperiodic lattices provide a controlled setting in which localization and spectral gaps arise without uncorrelated disorder. The Aubry-André-Harper (AAH) model serves as the paradigmatic example, featuring a self-dual localization transition AubryAndre1980 and is closely connected to the Harper equation and the Hofstadter problem Harper1955 ; Hofstadter1976 ; Kohmoto1983_PRL . Under OBC, the bulk energy gaps of the AAH spectrum can host spectrally isolated in-gap modes that are localized near the system boundaries, with their existence and spatial profiles depending on the quasiperiodic phase offset Thouless1983_Pump ; Kraus2012_PRL ; Kraus2013_PRL . Experimental realizations of quasiperiodic localization and gap physics have been reported in photonic and cold-atom platforms Celi2014_PRL ; Lahini2009_PRL ; Verbin2013_PRL ; Roati2008_Nature ; Schreiber2015_Science . Non-Hermitian extensions of AAH-type models, incorporating nonreciprocal hopping, complex quasiperiodic potentials, or pairing terms, have been demonstrated to alter localization properties and induce anomalous boundary or skin modes Ganeshan2015_PRL ; Longhi_NHAAH_PhaseTransitions ; Zeng2020_NHAAH_Topo ; Cai2021_NHAAH_pwave ; Li2023_Entanglement_NHAAH ; LocalizationTopo_NHQuasi . In contrast, the response of in-gap boundary-localized modes to a spatially fluctuating imaginary gauge field with zero mean (corresponding to the erratic-gauge regime) remains largely unexplored.
@@
These developments leave a specific gap: while quasiperiodic AAH chains can host spectrally isolated in-gap modes localized near open boundaries. However, the behavior of such boundary-localized in-gap modes remains unexplored when the hopping asymmetry arises from a zero-mean but strongly fluctuating imaginary gauge field-termed the erratic-gauge regime. In this work, we address this problem by employing an exact site-dependent nonunitary gauge transformation under OBC. This transformation maps the non-Hermitian chain onto its Hermitian AAH counterpart, preserving the open-boundary spectrum. The mapping shows that gauge fluctuations do not shift eigenvalues but reshape right/left eigenfunctions through a realization-dependent random-walk envelope. Within spectral gaps we identify two isolated boundary-localized modes with sharply different sensitivity to this envelope: one remains pinned to the boundary, whereas the other is gauge-steerable, allowing its probability maximum to be repositioned solely by varying the gauge realization without altering its eigenvalue. Furthermore, we propose a practical preparation protocol leveraging weak local gain. The gain site is determined once from the profile of the corresponding Hermitian mode and can then be fixed across different gauge realizations to selectively amplify the steerable in-gap mode from an initial bulk wave packet.
@@
The remainder of the paper is organized as follows. In Sec. II, the continuum and tight-binding models with Bernoulli gauge disorder is introduced. In Sec. III, the nonunitary gauge transformation is derived under OBC and used to analyze the Lyapunov exponents and the phase diagram is presented. In Sec. IV, we show the coexistence of pinned and steerable in-gap modes and verify this by finite-size diagnostics. In Sec. V, we propose a weak local-gain protocol and demonstrate the dynamical preparation of the gauge-steerable mode. Finally, we conclude in Sec. VI.
@@ II Model
$H_{\rm cont}=\frac{\bigl(p-i\hbar A(x)\bigr)^{2}}{2m}+V_{1}\sin^{2}(k_{1}x)+V_{2}\cos\!\bigl(2k_{2}x+\varphi\bigr),$ (1)
where $p=-i\hbar\partial_{x}$ is the momentum operator, $A(x)\in\mathbb{R}$ is a prescribed spatial profile (with dimension of inverse length), $V_{1,2}$ are the lattice depths, and $k_{1,2}$ are the wave numbers of the primary and secondary lattices, respectively. The phase $\varphi$ sets the spatial offset of the secondary lattice. We consider an open interval of space, introducing $X(x)=\int_{x_{0}}^{x}A(x^{\prime})\,dx^{\prime},$ so that $X^{\prime}(x)=A(x)$. Using $e^{-X}pe^{X}=p-i\hbar X^{\prime}(x)$, one obtains $e^{X}(p-i\hbar A)e^{-X}=p$ and hence the nonunitary gauge-transformed Hamiltonian reads
@@
In the deep-lattice regime $V_{1}\gg E_{R}$ (with recoil energy $E_{R}=\hbar^{2}k_{1}^{2}/2m$) and for a weaker secondary lattice $V_{1}\gg V_{2}$, we project $H_{\rm cont}^{H}$ to the lowest Bloch band of the primary lattice and retain only onsite and nearest-neighbor matrix elements. Let $a\equiv\pi/k_{1}$ be the primary-lattice spacing and $x_{n}\equiv na$ the lattice sites. Expanding the field operator in lowest-band Wannier functions $w(x-x_{n})$, the resulting tight-binding Hamiltonian takes the Hermitian AAH form
@@
Figure 1: Schematic of the AAH model with disordered imaginary gauge field. Blue (red) arrows indicate $J_{R}(n)$ $J_{L}(n)$]. The gray bars depict the onsite potential $V_{n}$ on each site, and the dashed curve is a guide to the quasiperiodic modulation profile.
@@
Finally, within the tight-binding description, we restore the imaginary gauge field of the original continuum model in Eq. (1) by applying the inverse nonunitary gauge transformation to $\hat{H}^{H}$. With $X_{n}\equiv X(x_{n})$, we define the site-diagonal operator $S=\exp\!\bigl(\sum_{n=1}^{N}X_{n}\hat{c}_{n}^{\dagger}\hat{c}_{n}\bigr)$ and obtain the non-Hermitian tight-binding Hamiltonian $\hat{H}=S\hat{H}^{H}S^{-1}$, i.e.,
$\hat{H}=\sum_{n=1}^{N}V_{n}\,\hat{c}_{n}^{\dagger}\hat{c}_{n}+\sum_{n=1}^{N-1}\!\left[J_{R}(n)\,\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+J_{L}(n)\,\hat{c}_{n}^{\dagger}\hat{c}_{n+1}\right].$ (5)
The right/left hopping amplitudes between sites $n$ and $n+1$ are
@@
where $s_{n}=\pm 1$ is a zero-mean Bernoulli sequence and $\Delta$ controls the disorder strength ENHSE_PRL2025 . The structure of the non-Hermitian AAH chain considered here is schematically illustrated in Fig. 1.
@@ III.1 Nonunitary gauge transformation
In the single-particle scenario, an eigenstate of Hamiltonian (5) can be written as $|\psi\rangle=\sum_{n=1}^{N}\psi_{n}\hat{c}_{n}^{\dagger}|0\rangle$. The coefficients $\psi_{n}$ and the corresponding eigenenergy $E$ satisfy the stationary Schrödinger equation $\hat{H}|\psi\rangle=E|\psi\rangle$, which in the site basis gives
@@
Thus the entire OBC spectrum of $\hat{H}$ coincides with that of the Hermitian AAH chain. To characterize the asymptotic spatial growth/decay of a given eigenstate in the thermodynamic limit, we introduce the Lyapunov exponent (spatial exponent)
$\Gamma(E)=\lim_{n\to\infty}\frac{1}{n}\ln\Big|\frac{\psi_{n}}{\psi_{1}}\Big|.$ (12)
@@
is the spatial average of the imaginary gauge field. For the Bernoulli gauge disorder defined in Sec. II, $\bar{h}=0$ for typical realizations in the long-chain limit, so that
@@
which yields the bulk localization length $\xi=1/\Gamma$ for $\lambda>2J$. For eigenvalues lying inside bulk gaps (e.g., edge modes under OBC), exponentially decaying solutions exist and their decay rates are generally energy dependent. Nevertheless, for any fixed eigenvalue $E$ the nonunitary gauge transformation implies that the spatial exponent extracted from $\psi_{n}$ differs from that of $\phi_{n}$ only through $\bar{h}$; hence for typical zero-mean gauge realizations with $\bar{h}=0$, the decay exponent of an in-gap state is inherited from the Hermitian counterpart.
@@ III.2 ENHSE-induced collective pinning and phase diagram
Figure 2: Static properties and phase structure for fixed $\Delta=0.4$, $\beta=(\sqrt{5}-1)/2$ and $\varphi=\pi/5$ hereafter for calculations. (a) Phase diagram in the $(J,\lambda)$ plane, in which $N=1000$, color coded by $S_{I}$. The dashed line $\lambda_{c}=2J$ marks the self-dual transition. The blue star and green circle indicate the parameter sets $(J,\lambda)=(1,1)$ and $(1,3)$, respectively. (b) Cumulative height $X_{n}$ (black, left axis) and averaged intensity $I_{n}$ (colored curves, right axis) for the two representative parameter sets marked in (a). (c),(d) Eigenstate density in the ENHSE regime $(J,\lambda)=(1,1)$ and in the AAH-localized regime $(J,\lambda)=(1,3)$, respectively.
To quantify the collective real-space structure of the spectrum of the non-Hermitian Hamiltonian (5), we introduce a spectrum-averaged intensity profile. Denoting the $N$ eigenstates by $\{\psi_{\alpha,n}\}$ with eigenstate index $\alpha=1,\dots,N$, we define
@@
which approaches its minimum value $S_{I}\simeq 1/N$ for a nearly uniform $I_{n}$, and increases when the spectral weight concentrates near a few preferred sites. Figure 2(a) shows the resulting phase diagram in the $(J,\lambda)$ plane (with $N=1000$), where the color encodes $S_{I}$. The dashed line $\lambda_{c}=2J$ marks the self-dual transition inherited from the Hermitian AAH counterpart, across which the Lyapunov exponent in Eq. (16) changes from $\Gamma=0$ to $\Gamma>0$. In the regime $\lambda<2J$, the Hermitian counterpart supports extended bulk eigenstates, yet the nonunitary envelope generated by the erratic imaginary gauge field can collectively concentrate the physical right eigenstates $\psi_{\alpha,n}$, yielding a strongly nonuniform $I_{n}$ and hence a large $S_{I}$. In contrast, for $\lambda>2J$ the quasiperiodic potential already localizes the eigenstates of the Hermitian counterpart, and the spectrum-averaged intensity becomes much less sensitive to the random-walk envelope, resulting in a significantly smaller $S_{I}$.
The microscopic origin of these distinct behaviors is illustrated in Fig. 2(b)–(d) for the two representative parameter sets marked in Fig. 2(a): $(J,\lambda)=(1,1)$ (blue star, $\lambda<2J$) and $(J,\lambda)=(1,3)$ (green circle, $\lambda>2J$). Figure 2(b) directly compares, for each parameter set, the cumulative random-walk profile $X_{n}$ (black curve, left axis) and the spectrum-averaged intensity $I_{n}$ defined in Eq. (17) (colored curve, right axis). For $(J,\lambda)=(1,1)$, where the eigenstates of the Hermitian counterpart are extended, the nonunitary envelope $e^{X_{n}}$ selects only a few preferred sites; accordingly, $I_{n}$ in Fig. 2(b) develops sharp peaks located near the dominant maxima of $X_{n}$. This collective pinning is further corroborated by the eigenstate-density shown in Fig. 2(c), where a large fraction of the spectrum accumulates its weight around the same spatial locations. For $(J,\lambda)=(1,3)$, by contrast, the Hermitian counterpart is in the AAH-localized regime, so individual eigenstates are exponentially localized with centers distributed across the chain. As a result, after averaging over the whole spectrum, $I_{n}$ in Fig. 2(b) does not display pronounced peaks but stays close to a weakly modulated background, showing no systematic correlation with the extrema of $X_{n}$. The eigenstate-density in Fig. 2(d) directly confirms that the states remain exponentially localized with distinct localization centers, consistent with the much smaller value of $S_{I}$ in this regime.

### v2

@@ Abstract
We identify steerable exponentially localized in-gap mode in a quasiperiodic non-Hermitian Aubry-André-Harper chain with a spatially fluctuating, zero-mean imaginary gauge field. Under open boundary conditions, the system is exactly related to the Hermitian AAH model by a nonunitary gauge transformation: the OBC spectrum and Lyapunov exponents are unchanged, while eigenstates acquire a gauge-dependent envelope. In a parameter region with spectrally isolated in-gap boundary modes, we find two exponentially localized in-gap modes with sharply different responses to the imaginary gauge field. One remains boundary pinned, but the other is gauge-steerable: it stays exponentially localized while its probability maximum shifts as the gauge field is changed, with its eigenenergy unchanged. We further show that weak on-site gain, applied at a single site chosen once and then kept fixed, can dynamically prepare this steerable mode from a generic bulk wave packet. Changing the gauge field then yields exponentially localized states at different locations.
@@ I Introduction
Non-Hermitian Hamiltonians provide an effective description of driven or open quantum systems with gain and loss, leakage, measurement backaction, or nonreciprocal couplings. Such systems host phenomena absent in Hermitian settings, including parity-time symmetry breaking, exceptional points, and complex-spectral topology Bender1998_PRL ; Bender2007_RPP ; Rotter2009_JPA ; Heiss2012_JPA ; ElGanainy2018_NatPhys ; MiriAlu2019_Science ; Ozdemir2019_NatMat ; Ruter2010_NatPhys ; Regensburger2012_Nature ; Hodaei2017_Nature ; Feng2017_NatPhoton ; Bergholtz2021_RMP ; Ashida2020_AdvPhys . In lattice systems, non-Hermiticity can strongly reshape localization and boundary responses, enabling boundary-state accumulation, amplification, and wave manipulation beyond Hermitian paradigms NHSE_review_Zhang2022 ; NonHermitianReview_Kawabata2020 ; 1Manna20236 .
@@
A paradigm of boundary sensitivity is the non-Hermitian skin effect (NHSE). Under open boundary conditions (OBC), a macroscopic number of eigenstates can become exponentially localized near a boundary, in stark contrast to the Bloch-like states of the same model under periodic boundary conditions (PBC). This breakdown of conventional Bloch band theory has motivated generalized bulk–boundary correspondences based on non-Bloch band theory and point-gap topology YaoWang2018_PRL ; Kunst2018_PRL ; Gong2018_PRX ; YokomizoMurakami2019_PRL ; Kawabata2019_PRX ; Okuma2020_PRL ; Lee2016_PRL ; Borgnia2020_PRL124056802 , and the resulting boundary spectra and signatures have been extensively investigated Bergholtz2021_RMP ; Ashida2020_AdvPhys ; NHSE_review_Zhang2022 ; NonHermitianReview_Kawabata2020 .
@@
Beyond uniform nonreciprocity, imaginary gauge fields offer a distinct route to non-Hermitian localization. In the Hatano-Nelson model, a uniform imaginary vector potential modifies localization and induces strong boundary sensitivity HatanoNelson1996_PRL ; HatanoNelson1997_PRB ; PRL119137402 ; PRL119137402 ; PRB103064201 . More recently, Longhi showed that skin-like localization can arise purely from spatial fluctuations of the imaginary gauge field even when its spatial average vanishes, so that the lattice remains globally reciprocal ENHSE_PRL2025 . In this erratic non-Hermitian skin effect (ENHSE), eigenstates acquire an effective random envelope, and their localization centers are controlled by extrema of the cumulative gauge field. This mechanism differs from the conventional NHSE driven by net nonreciprocity and from Anderson localization. A natural next question is how this envelope reshaping affects spectrally isolated in-gap boundary modes that are already exponentially localized.
@@
A natural arena for addressing this question is provided by quasiperiodic lattices, where localization and spectral gaps emerge without uncorrelated disorder. The Aubry-André-Harper (AAH) model is a canonical example, featuring a self-dual localization transition AubryAndre1980 and close connections to the Harper equation and the Hofstadter problem Harper1955 ; Hofstadter1976 ; Kohmoto1983_PRL . Under OBC, bulk gaps can host spectrally isolated in-gap modes localized near system boundaries, with their existence and profiles depending on the quasiperiodic phase offset Thouless1983_Pump ; Kraus2012_PRL ; Kraus2013_PRL . Quasiperiodic localization and gap physics have been realized in photonic and cold-atom platforms Celi2014_PRL ; Lahini2009_PRL ; Verbin2013_PRL ; Roati2008_Nature ; Schreiber2015_Science . Non-Hermitian extensions of AAH-type models have also been studied, showing modified localization and anomalous boundary or skin modes Ganeshan2015_PRL ; Longhi_NHAAH_PhaseTransitions ; Zeng2020_NHAAH_Topo ; Cai2021_NHAAH_pwave ; Li2023_Entanglement_NHAAH ; LocalizationTopo_NHQuasi . What remains unclear is how these spectrally isolated in-gap boundary modes respond when the hopping asymmetry is generated by a zero-mean but strongly fluctuating imaginary gauge field (the erratic-gauge regime).
@@
In this work, we address this question by employing an exact nonunitary gauge transformation under OBC. This transformation maps the non-Hermitian chain onto its Hermitian AAH counterpart, so the open-boundary spectrum and Lyapunov exponents are unchanged, while right/left eigenfunctions are reshaped by a spatial envelope set by the gauge field. Within spectral gaps, we identify two isolated in-gap modes with sharply different responses to the gauge field: one remains pinned to the boundary, whereas the other stays exponentially localized but becomes gauge steerable, with its probability peak repositioned by varying the gauge field without shifting its eigenvalue. We further propose a preparation protocol based on weak local gain: the gain is applied at a single site selected from the Hermitian counterpart mode profile and then kept fixed, which allows the steerable in-gap mode to be amplified from a bulk wave packet and, by changing the gauge field, to be prepared at different locations.
@@
The remainder of the paper is organized as follows. In Sec. II, the continuum and tight-binding models with Bernoulli gauge disorder is introduced. In Sec. III, the nonunitary gauge transformation is derived under OBC and used to analyze the Lyapunov exponents and the phase diagram is presented. In Sec. IV, we show the coexistence of pinned and steerable in-gap modes and verify this by finite-size adjudication. In Sec. V, we propose a weak local-gain protocol and demonstrate the dynamical preparation of the gauge-steerable mode. Finally, we conclude in Sec. VI.
@@ II Model
$H_{\rm cont}=\frac{\bigl(p-iA(x)\bigr)^{2}}{2m}+V_{1}\sin^{2}(k_{1}x)+V_{2}\cos\!\bigl(2k_{2}x+\varphi\bigr),$ (1)
where $p=-i\partial_{x}$ is the momentum operator ($\hbar=1$, hereafter), $A(x)\in\mathbb{R}$ is a random non-Hermitian external field HatanoNelson1996_PRL ; HatanoNelson1997_PRB ; PRL119137402 ; PRL119137402 ; PRB103064201 , $V_{1,2}$ are the lattice depths, and $k_{1,2}$ are the wave numbers of the primary and secondary lattices, respectively. The phase $\varphi$ sets the spatial offset of the secondary lattice. We focus on a chain with open boundaries, introducing $X(x)=\int_{x_{0}}^{x}A(x^{\prime})\,dx^{\prime},$ so that $X^{\prime}(x)=A(x)$. Using $e^{-X}pe^{X}=p-iX^{\prime}(x)$, one obtains $e^{X}(p-iA)e^{-X}=p$ and hence the nonunitary gauge-transformed Hamiltonian reads
@@
In the deep-lattice regime $V_{1}\gg E_{R}$ (with recoil energy $E_{R}=k_{1}^{2}/2m$) and for a weaker secondary lattice $V_{1}\gg V_{2}$, we project $H_{\rm cont}^{H}$ to the lowest Bloch band of the primary lattice and retain only onsite and nearest-neighbor matrix elements. Let $a\equiv\pi/k_{1}$ be the primary-lattice spacing and $x_{n}\equiv na$ the lattice sites. Expanding the field operator in lowest-band Wannier functions $w(x-x_{n})$, the resulting tight-binding Hamiltonian takes the Hermitian AAH form
@@
Figure 1: Schematic of the AAH model with disordered imaginary gauge field. Blue (red) arrows indicate $J_{R}(n)$ $J_{L}(n)$]. The gray bars depict the onsite potential $V_{n}$ on each site, and the dashed curve is a guide for the quasiperiodic modulation profile.
@@
Finally, within the tight-binding description, we restore the imaginary gauge field of the original continuum model in Eq. (1) by applying the inverse nonunitary gauge transformation to $\hat{H}^{H}$ with $X_{n}\equiv X(x_{n})$. We define the transformation operator $S=\exp\!\bigl(\sum_{n=1}^{N}X_{n}\hat{c}_{n}^{\dagger}\hat{c}_{n}\bigr)$ and obtain the non-Hermitian tight-binding Hamiltonian by the nonunitary transformation,
$\hat{H}$ $=S\hat{H}^{H}S^{-1}$
$=\sum_{n=1}^{N}V_{n}\,\hat{c}_{n}^{\dagger}\hat{c}_{n}+\sum_{n=1}^{N-1}\!\left[J_{R}(n)\,\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+J_{L}(n)\,\hat{c}_{n}^{\dagger}\hat{c}_{n+1}\right].$ (5)
The right/left hopping amplitudes are
@@
where $\{s_{n}\}=\pm 1$ is a zero-mean Bernoulli sequence and $\Delta$ controls the disorder strength ENHSE_PRL2025 . This means the random non-Hermitian external field obeys Bernoulli distribution. The structure of the non-Hermitian AAH chain considered here is schematically illustrated in Fig. 1.
@@ III.1 Nonunitary gauge transformation
To obtain an exact and analytically tractable description of the phase structure and in-gap modes in the presence of an erratic imaginary gauge field, we map the non-Hermitian Hamiltonian under OBC to its Hermitian AAH counterpart via a nonunitary gauge transformation, so that eigenvalues are preserved and non-Hermitian effects enter only through the spatial rescaling of eigenstates. In the single-particle scenario, an eigenstate of Hamiltonian (5) can be written as $|\psi\rangle=\sum_{n=1}^{N}\psi_{n}\hat{c}_{n}^{\dagger}|0\rangle$. The coefficients $\psi_{n}$ and the corresponding eigenenergy $E$ satisfy the stationary Schrödinger equation $\hat{H}|\psi\rangle=E|\psi\rangle$, which in the site basis gives
@@
Thus the entire OBC spectrum of $\hat{H}$ coincides with the Hermitian counterpart AAH chain. To characterize the asymptotic spatial growth/decay of a given eigenstate in the thermodynamic limit, we introduce the Lyapunov exponent (spatial exponent)
$\Gamma(E)=\left|\lim_{n\to\infty}\frac{1}{n}\ln\left|\frac{\psi_{n}}{\psi_{1}}\right|\right|.$ (12)
@@
is the spatial average of the imaginary gauge field. For the Bernoulli gauge disorder defined by Eq. 8, $\bar{h}=0$ in the long-chain limit, so that
@@
which yields the bulk localization length $\xi=1/\Gamma(E)$ for $\lambda>2J$.
For OBC eigenvalues $E$ that fall inside a bulk gap, the associated transfer matrix has a nonzero Lyapunov exponent, implying an exponentially decaying tail whose decay rate is energy dependent AubryAndre1980 ; Kohmoto1983_PRL ; Thouless1983_Pump . Nevertheless, for any fixed eigenvalue $E$ the nonunitary gauge transformation implies that the spatial exponent extracted from $\psi_{n}$ differs from that of $\phi_{n}$ only through $\bar{h}$; hence for typical zero-mean gauge fields with $\bar{h}=0$, the decay exponent of an in-gap state is inherited from the Hermitian counterpart.
@@ III.2 ENHSE-induced collective pinning and phase diagram
Figure 2: Phase structure and distribution of eigenstates when $\Delta=0.4$, $\beta=(\sqrt{5}-1)/2$ and $\varphi=\pi/5$ hereafter for calculations. (a) Phase diagram indicated by $S_{I}$ in Eq. 18 versus $J$ and $\lambda$ when $N=1000$. $\lambda_{c}=2J$ is the demarcation line for different phases. The blue star and green circle indicate the parameter sets $(J,\lambda)=(1,1)$ and $(1,3)$, respectively. (b) $X_{n}$ defined in Eq. 7 (black, left axis) and $I_{n}$ defined in Eq. 17 (colored curves, right axis) for the two parameter sets marked in (a). (c), (d) Eigenstate density in the ENHSE regime $(J,\lambda)=(1,1)$ and in the AAH-localized regime $(J,\lambda)=(1,3)$, respectively.
To quantify the collective real-space structure of the spectrum of the non-Hermitian Hamiltonian (5), we introduce a statistical quantity for a site $n$ involving all eigenstates. Denoting the $N$ eigenstates by $\{\psi_{\alpha,n}\}$ with eigenstate index $\alpha=1,\dots,N$, we define
@@
which approaches its minimum value $S_{I}\simeq 1/N$ for a nearly uniform $I_{n}$, and increases when major of the eigenstates with distribution concentrates near a few preferred sites. Figure 2(a) shows the resulting phase diagram indicated by $S_{I}$ in the $(J,\lambda)$ plane (with $N=1000$). The dashed line $\lambda_{c}=2J$ marks the self-dual transition inherited from the Hermitian AAH counterpart, across which the Lyapunov exponent in Eq. (16) changes from $\Gamma=0$ to $\Gamma>0$. In the regime $\lambda<2J$, the Hermitian counterpart supports extended bulk eigenstates, yet the nonunitary envelope generated by the erratic imaginary gauge field can collectively concentrate the physical right eigenstates $\psi_{\alpha,n}$, yielding a strongly nonuniform $I_{n}$ and hence a large $S_{I}$. In contrast, for $\lambda>2J$ the quasiperiodic potential already localizes the eigenstates of the Hermitian counterpart, and the spectrum-averaged intensity becomes much less sensitive to the envelope of $X_{n}$, resulting in a significantly smaller $S_{I}$.
The statistical details of the distinct behavior are illustrated in Fig. 2(b)-(d) for the two representative parameter sets marked in Fig. 2(a): $(J,\lambda)=(1,1)$ (blue star, $\lambda<2J$) and $(J,\lambda)=(1,3)$ (green circle, $\lambda>2J$). Figure 2(b) directly compares, for each parameter set, the cumulative profile $X_{n}$ (black curve, left axis) and $I_{n}$ defined in Eq. (17) (colored curve, right axis). For $(J,\lambda)=(1,1)$, where the eigenstates of the Hermitian counterpart are extended, the nonunitary envelope $e^{X_{n}}$ selects only a few preferred sites; accordingly, $I_{n}$ in Fig. 2(b) develops sharp peaks located near the dominant maxima of $X_{n}$. This collective pinning is further corroborated by the eigenstate-density shown in Fig. 2(c), where a large fraction of the eigenstates accumulates its weight around the same spatial locations. For $(J,\lambda)=(1,3)$, by contrast, the Hermitian counterpart is in the AAH-localized regime, so individual eigenstates are exponentially localized with centers distributed across the chain. As a result, $I_{n}$ in Fig. 2(b) does not display pronounced peaks but stays close to a weakly modulated background, showing no systematic correlation with the extreme of $X_{n}$. The eigenstate-density in Fig. 2(d) confirms that the states remain exponentially localized with distinct localization centers, consistent with the much smaller value of $S_{I}$ in this region.